WACC AND COST OF COMMON EQUITY Kahn Inc. has a target capital structure of 55% common equity and 45% debt to fund its $11 bilI will rate!
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A B C D 1 2. 3 4. 5 Debt ratio Equity ratio Total operating assets Debt Equity 45% 55% $11.00 billion $4.95 billion $6.05 bil

Cell reference -

A B с D 1 0.45 2 3 0.55 4 Debt ratio Equity ratio Total operating assets Debt Equity 11 =C4*C2 =C4*C3 billion billion billion

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.